Jeff Krause purchased 1000 shares of speculative stock on January 2 for 2.07 per share. Six months later on July 1 he sold them for 10.06 per share he uses a online broker that charges him 10 per a trade what was Jeff’s annualized HPR on this investment
Answer #1

Sale price=10.06 per share,

So total proceeds before brokerage = (1000*10.06) = 10,060

  Less: brockerage = 10,

Net proceeds =    10,050

Less: cost of Purchase =1000*2.007 = 2,070

Profit for the holding period( Six Months= 7,980

% for 6 months = 7980/2010*100 = 385.507246376812%, annualisede rate =   385.507246376812% *(12/6) =385.51% PA